OpenStudy (anonymous):

solve word problems that involve linear equations. Jack works at a bike store. Jack earns $150 every week plus $25 for every bike (b) that he sells. If Jack receives his paycheck at the end of the week for $325, how many bikes did he sell? Equation:__________ Answer:____________

OpenStudy (anonymous):

150 + 25b = 325 (b represents the number of bikes sold) Now solve for b. Can you take it from here ?

OpenStudy (anonymous):

b=7

OpenStudy (anonymous):

let me check... 150 + 25b = 325 25b = 325 - 150 25b = 175 b = 175/25 b = 7 You are correct. Jack sold 7 bikes. Good job :)

OpenStudy (anonymous):

thanks kellie

OpenStudy (anonymous):

your welcome....I am glad to help :)